# Optimal Sparsity of Mixture-of-Experts Language Models for Reasoning Tasks

### v1

@@ Abstract
Empirical scaling laws have driven the evolution of large language models (LLMs), yet their coefficients shift whenever the model architecture or data pipeline changes. Mixture‑of‑Experts (MoE) models, now standard in state‑of‑the‑art systems, introduce a new sparsity dimension that current dense‑model frontiers overlook. We investigate how MoE sparsity influences two distinct capability regimes: memorization and reasoning. We train families of MoE Transformers that systematically vary total parameters, active parameters, and top‑$k$ routing while holding the compute budget fixed. For every model we record pre‑training loss, downstream task loss, and task accuracy, allowing us to separate the train‑test generalization gap from the loss‑accuracy gap. Memorization benchmarks improve monotonically with total parameters, mirroring training loss. By contrast, reasoning performance saturates and can even regress despite continued gains in both total parameters and training loss. Altering top‑$k$ alone has little effect when active parameters are constant, and classic hyperparameters such as learning rate and initialization modulate the generalization gap in the same direction as sparsity. Neither post‑training reinforcement learning (GRPO) nor extra test‑time compute rescues the reasoning deficit of overly sparse models. Our model checkpoints, code and logs are open‑source at https://github.com/rioyokotalab/optimal-sparsity.
Machine Learning, ICML
@@ 1 Introduction
The recent evolution of large language models (LLMs) has been driven by empirical scaling laws (Hestness et al., 2017) that link training loss to model size, dataset size, and compute budget. Kaplan et al. showed that these laws hold across seven orders of magnitude, establishing them as a reliable extrapolation tool for dense Transformers (Kaplan et al., 2020). Subsequent work by Hoffmann et al. demonstrated that scaling curves can be inverted to choose the compute-optimal combination of parameters and tokens for a fixed budget (Hoffmann et al., 2022). Together, these results have made scaling analysis a cornerstone of model planning at both academic and industrial labs.
@@
Yet the coefficients of the scaling laws are not universal. Highly expressive models trained under different optimizers or architectures often follow the same loss trajectory but diverge substantially on downstream reasoning benchmarks (Liu et al., 2023). Brandfonbrener et al. extend the classic laws with loss-to-loss prediction, showing that the mapping between training and test distributions admits its own power law when the distributions differ substantially (Brandfonbrener et al., 2025). These observations imply that optimal budgets must be re-estimated whenever we modify the model or the data pipeline.
@@
A particularly compelling architectural modification is the Mixture-of-Experts (MoE) paradigm, offering high capacity at fixed FLOPs by routing each token through a sparse subset of experts (Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2021). Modern flagship models, e.g., Gemini 2.5 Pro (Gemini Team, 2025), DeepSeek-V3 (DeepSeek-AI, 2025b), and Qwen3 (Qwen Team, 2025) now rely on MoE as a de-facto standard for economical scaling. Abnar et al. derive a parameters-vs-FLOPs frontier and locate an optimal sparsity for a given compute budget (Abnar et al., 2025). These findings emphasize that the classical dense-model frontier is an incomplete picture, and one must account for architectural knobs such as MoE sparsity and top-$k$ routing.
@@
Furthermore, loss-based scaling curves do not always predict the performance on downstream tasks. Jelassi et al. report that increasing MoE sparsity improves memorization benchmarks, but saturates for reasoning performance (Jelassi et al., 2025). However, the Mixture of Parrots paper (Jelassi et al., 2025) only explores the number of active vs. total parameters, ignoring the effect of routing strategies beyond standard top-2 routing. They also do not consider the effect of reinforcement learning and test-time compute on their reasoning benchmarks. Evaluating reasoning performance immediately after pre-training overlooks both the benefits of post-training adaptation and the leverage of additional test-time compute. Post-training methods such as GRPO, which use reinforcement signals to encourage coherent chain-of-thought generation, sharpen a model’s reasoning on complex tasks (OpenAI, 2024b; DeepSeek-AI, 2025a). Beyond these refinements, models can further improve outputs at test time by adopting calibrated decoding strategies that mirror how humans pause to reconsider difficult problems. These test-time approaches not only boost routine benchmark performance but, when properly tuned, substantially enhance multi-step mathematical reasoning, demonstrating that adaptive computing at test time is a powerful complement to both model scale and post-training adaptation.
@@
In this paper, we aim to identify how the optimal sparsity of MoE changes between memorization (TriviaQA, HellaSwag) and reasoning (GSM8K, GSM-Plus) tasks. In this work, we use the term dense models to refer to standard Transformers with a single feed-forward network per layer. For MoE models, we define sparsity as $\text{sparsity}=1-\tfrac{\text{Top-}k}{\text{Experts}}$ following the convention that sparsity measures the fraction of inactive parameters. We train families of MoEs varying not only the total vs. active parameters, but also the number of top-$k$ experts. For each model, we measure the loss on the pre-training data, the task loss on the downstream benchmarks, and the accuracy on those benchmarks. This allows us to disentangle the generalization gap between the train vs. test loss, and the gap between loss vs. accuracy. For both memorization and reasoning benchmarks, the train loss decreases monotonically with the total parameters. The task loss and accuracy follow the same monotonic trend as the train loss for memorization benchmarks. In contrast, for reasoning benchmarks, the task loss and accuracy diverge from the monotonic trend as the total parameters increase and training loss decreases. We found that changing the $k$ in top-$k$ routing itself has a negligible effect if the number of active parameters is kept constant. We also consider classic generalization-gap controls by sweeping the learning rate and initialization, and show that their effects align strikingly with the generalization-gap caused by sparsity. This confirms that the gap between the performance on memorization vs. reasoning tasks can be induced not only by sparsity of the MoE, but also classical hyperparameters like learning rate and initialization. We further investigate whether applying GRPO or additional test-time compute could recover the poor reasoning ability of sparser models. Our results show that the gap between memorization and reasoning performance caused by increased sparsity remains unchanged even after GRPO and increased test-time compute. This means that finding the optimal sparsity of the MoE during pre-training is crucial for training a reasoning model under a fixed compute budget. We release model checkpoints, code and logs are open-source at https://github.com/rioyokotalab/optimal-sparsity.
@@ 2.1 Mixture of Experts
Mixture-of-Experts (MoE) networks were introduced by (Jacobs et al., 1991; Jordan & Jacobs, 1994) and later brought to large-scale neural language modeling by Shazeer et al. (2017). Within the Transformer architecture (Vaswani et al., 2017), MoE layers have proven especially effective, scaling to hundreds of billions of parameters while maintaining manageable training costs (Lepikhin et al., 2021; Fedus et al., 2021; Du et al., 2022; Zoph et al., 2022). Consequently, modern state-of-the-art language models, including Gemini 2.5 Pro (Gemini Team, 2025), DeepSeek-V3 (DeepSeek-AI, 2025b), and Qwen3 (Qwen Team, 2025), rely heavily on MoE layers to achieve superior performance under fixed inference budgets. In an MoE layer, a learnable router assigns each token to a sparse subset of experts. Let $\mathbf{x}\in\mathbb{R}^{d_{h}}$ be a token representation and $\{\text{FFN}(\mathbf{x})_{i}\}_{i=1}^{n}$ the $n$ feed-forward experts. For top-$k$ routing, the router produces scores $\mathbf{s}=\mathbf{x}^{\top}\mathbf{W}_{\text{router}}\in\mathbb{R}^{n},$ and selects the indices $\mathcal{K}$ of the $k$ largest components, then normalizes them: $g(\mathbf{x})_{i}=\frac{\exp(s_{i})}{\sum_{j\in\mathcal{K}}\exp(s_{j})}$ if $i\in\mathcal{K}$ and $g(\mathbf{x})_{i}=0$ otherwise.
@@ 2.2 Scaling Laws of LLMs
Existing scaling laws demonstrate power-law relationships between model performance, parameter count, dataset size, and compute budget (Kaplan et al., 2020; Hoffmann et al., 2022). Scaling laws for MoE models have similarly explored how total parameter count and expert granularity jointly affect scaling behavior (Clark et al., 2022; Ludziejewski et al., 2024). Building on this, Frantar et al. derived sparsity-aware scaling exponents that bridge dense and sparse regimes (Frantar et al., 2024), while Abnar et al. empirically charted the optimal trade-offs between total parameters and FLOPs per token in MoE settings (Abnar et al., 2025). Complementary theoretical and empirical work shows that adding experts tends to improve memorization more than reasoning, motivating new, generalized scaling frameworks that address scaling laws for reasoning performance (Jelassi et al., 2025). Furthermore, recent analyses indicate that increasing sparsity itself can directly improve loss, highlighting sparsity as a key dimension in scaling behavior (Team, 2025).
@@
Since the scaling law for next-token prediction loss does not necessarily align with downstream task loss, it may not be reliable for predicting benchmark performance (Grattafiori et al., 2024). Some work has tried to model downstream accuracy with an exponential curve, but accuracy is only predictable when we average over many tasks and carefully choose which ones to include (Gadre et al., 2024). Another line of research instead first quantifies how downstream task loss scales with parameters and data, then converts predicted losses into accuracy estimates, achieving under two points of absolute error for mid-scale models using minimal extra compute (Bhagia et al., 2024). Prior work observes that downstream task loss relates to pre-training loss, where the shifts depend on the minimal achievable losses determined by the intrinsic complexity and distributional mismatch between the pre-training and downstream datasets (Brandfonbrener et al., 2025). Because scaling laws differ across tasks, the optimal scaling strategy may also vary; for example, knowledge-based QA tasks are “capacity-hungry,” benefiting more from larger model sizes, whereas code-related tasks are “data-hungry,” benefiting more from increased training data (Roberts et al., 2025).
Inverse Scaling.
Standard scaling laws imply that one can indefinitely increase model and dataset size without worrying about overfitting; however, if we need to worry about generalization and overfitting, these scaling laws break down (Caballero et al., 2023). Through the Inverse Scaling Prize, Wei et al. and McKenzie et al. systematically identified tasks where larger models perform worse (Wei et al., 2023; McKenzie et al., 2023). More recently, Lourie et al argue that scaling laws are unreliable predictors of downstream task performance, framing inverse scaling as a practical constraint on model development (Lourie et al., 2025).
@@ 3 Experiments
In this section, we empirically demonstrate the scaling of downstream task performance through a systematic investigation of memorization-reasoning benchmarks in MoE LLMs.
@@ 3.1 Experimental Setup
We use a balanced mixture of general-domain and mathematics-centric corpora, totaling 125 B-tokens. High quality web text (43 B) comes from de-duplicated DCLM (Zyphra, 2024), the Flan-decontaminated Dolmino subset, and WebInstructFull. Mathematics (32 B) combines OLMo OpenWebMath and Algebraic-Stack (Soldaini et al., 2024), FineMath-4+ (Liu et al., 2024a), the MathPile commercial subset (Wang et al., 2024), the math split of Dolmino-Mix-1124 (OLMo, 2025), OpenMathInstruct-1/2 (Toshniwal et al., 2024b, a), StackMathQA, Orca-Math (Mitra et al., 2024), and GSM8K train (Cobbe et al., 2021). STEM Literature & Reference (42 B) consists of arXiv, pes2o, Wikipedia, and Dolma–books (Soldaini et al., 2024). Finally, we add Code from the StackExchange code subset. See Appendix A.1 for complete statistics.
@@
We evaluate three capability areas with standard few-shot prompts. Mathematical Reasoning: GSM8K (Cobbe et al., 2021) (4-shot) and GSM-Plus (Li et al., 2024) (5-shot CoT). Reading Comprehension: TriviaQA (Joshi et al., 2017) with 4-shot prompting. Commonsense Reasoning: HellaSwag (Zellers et al., 2019), each under a 4-shot prompting setup. See Appendix 3 for further details.
@@ 3.2 Downstream Performance Does Not Necessarily Improve with Total Parameter Size
Figure 1 presents the training and validation losses when fixing the top-kk/MoE layer width constant and increasing only the number of experts (and hence the total parameter count). As the total parameter count grows, both training and validation losses decrease. Therefore, in terms of pre-training loss, increasing total parameters (thereby raising sparsity) reduces pre-training loss, which is consistent with prior work.
@@
Can we avoid a decline in performance as the total number of experts increases? Figure 2 shows that models with more active parameters begin to overfit at a lower pre-training loss and reach a lower minimum task loss at their optimal expert counts. Consequently, improving results on GSM8K and GSM-Plus requires tuning not only the total number of experts but also the top-$k$ size. Whether a similar trend occurs on other reasoning benchmarks, including code-generation tasks, remains an open question.
@@
The decline in math-task performance as total parameters increase is not limited to task loss; it also consistently holds for downstream accuracy (Figure 3). For TriviaQA and HellaSwag, accuracy improves monotonically as training loss decreases. By contrast, on GSM8K, further reductions in pre-training loss do not always translate to higher accuracy. When the number of active parameters is held constant, over-optimizing pre-training loss can indeed harm performance. Figure 4 plots benchmark error rate against pre-training loss, including intermediate checkpoints. We observe a sparsity dependence for reasoning-oriented tasks such as GSM8K and GSM-Plus. These results suggest that, for MoE models, downstream accuracy can deviate from the predictions of conventional scaling laws, and these deviations may vary across different tasks.
@@
Figure 4: Effect of sparsity on performance across different tasks We vary sparsity (1 - top-kk/Experts) and plot the relationship between pre-training loss and benchmark error rate, including intermediate checkpoints. For TriviaQA and HellaSwag, the error rate clearly tracks training loss and is largely insensitive to sparsity. In contrast, reasoning-intensive tasks such as GSM8K and GSM-Plus exhibit a strong dependence of error rate on sparsity.

### v2

@@ Abstract
Empirical scaling laws have driven the evolution of large language models (LLMs), yet their coefficients shift whenever the model architecture or data pipeline changes. Mixture-of-Experts (MoE) models, now standard in state-of-the-art systems, introduce a new sparsity dimension that current dense-model frontiers overlook. We investigate how MoE sparsity influences two distinct capability regimes: memorization skills and reasoning skills. By training MoE families that vary total parameters, active parameters, and top-$k$ routing under fixed compute budgets, we disentangle pre-training loss from downstream accuracy. Our results reveal two principles. First, Active FLOPs: models with identical training loss but greater active compute achieve higher reasoning accuracy. Second, Total tokens per parameter (TPP): memorization tasks improve with more parameters, while reasoning tasks benefit from optimal TPP, indicating that reasoning is data-hungry. Neither reinforcement learning post-training (GRPO) nor increased test-time compute alters these trends. We therefore argue that optimal MoE sparsity must be determined jointly by active FLOPs and TPP, revising the classical picture of compute-optimal scaling. Our model checkpoints, code and logs are open-source at https://github.com/rioyokotalab/optimal-sparsity.
@@ 1 Introduction
The recent evolution of large language models (LLMs) has been driven by empirical scaling laws (Hestness et al., 2017) that link training loss to model size, dataset size, and compute budget. Kaplan et al. (2020) showed that these laws hold across seven orders of magnitude, establishing them as a reliable extrapolation tool for dense Transformers. Subsequent work by  Hoffmann et al. (2022) demonstrated that scaling curves can be inverted to choose the compute-optimal combination of parameters and tokens for a fixed budget. Together, these results have made scaling analysis a cornerstone of model planning at both academic and industrial labs.
@@
Yet the coefficients of the scaling laws are not universal. Highly expressive models trained under different optimizers or architectures often follow the same loss trajectory but diverge substantially on downstream reasoning benchmarks (Liu et al., 2023). Brandfonbrener et al. (2025) extend the classic laws with loss-to-loss prediction, showing that the mapping between training and test distributions admits its own power law when the distributions differ substantially. These observations imply that optimal budgets must be re-estimated whenever we modify the model or the data pipeline.
@@
A particularly compelling architectural modification is the Mixture-of-Experts (MoE) paradigm, offering high capacity at fixed FLOPs by routing each token through a sparse subset of experts (Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2021). Modern flagship models, e.g., Gemini 2.5 Pro (Gemini Team, 2025), DeepSeek-V3 (DeepSeek-AI, 2025b), and Qwen3 (Qwen Team, 2025) now rely on MoE as a de facto standard for economical scaling. Abnar et al. (2025) derive a parameters-vs-FLOPs frontier and locate an optimal sparsity for a given compute budget. These findings emphasize that the classical dense-model frontier is an incomplete picture, and one must account for architectural knobs such as MoE sparsity and top-$k$ routing.
@@
Evaluating reasoning performance immediately after pre-training overlooks both the benefits of post-training adaptation and the leverage of additional test-time compute. Post-training methods such as GRPO, which use reinforcement signals to encourage coherent chain-of-thought generation, sharpen a model’s reasoning on complex tasks (OpenAI, 2024b; DeepSeek-AI, 2025a). Beyond these refinements, models can further improve outputs at test time by adopting calibrated decoding strategies that mirror how humans pause to reconsider difficult problems. These test-time approaches not only boost routine benchmark performance but, when properly tuned, substantially enhance multi-step mathematical reasoning, demonstrating that adaptive computing at test time is a powerful complement to both model scale and post-training adaptation.
@@
In this paper, we aim to identify how the optimal sparsity of MoE changes between memorization skills (TriviaQA, HellaSwag) and reasoning skills (GSM8K, GSM-Plus) tasks. In this work, we use the term dense models to refer to standard Transformers with a single feed-forward network per layer. For MoE models, we define sparsity as $\text{sparsity}=1-\tfrac{\text{Top-}k}{\text{Experts}}$ following the convention that sparsity measures the fraction of inactive parameters. We train families of MoEs varying not only the total vs. active parameters, but also the number of top-$k$ experts. For each model, we measure the loss on the pre-training data, the task loss on the downstream benchmarks, and the accuracy on those benchmarks. This allows us to disentangle the generalization gap between the train vs. test loss, and the gap between loss vs. accuracy. For both memorization and reasoning benchmarks, the train loss decreases monotonically with the total parameters count increases. The task loss and accuracy follow the same monotonic trend as the train loss for memorization benchmarks. In contrast, for reasoning benchmarks, the task loss and accuracy diverge from this monotonic trend as the total parameters increase and training loss decreases. We find that changing the $k$ in top-$k$ routing has a negligible effect if the number of active parameters is kept constant. We also consider classic generalization-gap controls by sweeping the learning rate and initialization, showing that their effects align strikingly with the generalization-gap caused by sparsity. This confirms that the gap between the performance on memorization skills and reasoning skills can be induced not only by the sparsity of the MoE, but also by classical hyperparameters like learning rate and initialization. We further investigate whether applying GRPO or additional test-time compute could recover the degraded reasoning ability of sparser models. Our results show that the gap between memorization and reasoning performance caused by increased sparsity remains unchanged even after GRPO and increased test-time compute. This means that finding the optimal sparsity of the MoE during pre-training is crucial for training a reasoning model under a fixed compute budget.
We characterize these divergences along two key axes: (1) Active FLOPs - downstream reasoning performance is not determined by training loss alone, but by the number of active FLOPs at both train and test time; even at identical training loss, models with a larger top-$k$ consistently outperform smaller ones. (2) Total tokens per parameter (TPP) - reasoning ability peaks around 20 tokens per parameter, whereas memorization skills are parameter-hungry and improve with lower TPP. Together, these axes define the compute-optimal sparsity for MoE models.
We further demonstrate that neither reinforcement learning post-training (GRPO) nor additional test-time compute eliminates this trade-off, highlighting that pre-training sparsity remains the dominant factor for reasoning ability under fixed budgets.
Together, our findings refine the scaling laws for MoE LLMs: memorization improves with higher sparsity and more experts whereas reasoning requires a careful balance between active FLOPs and data-per-parameter, occasionally favoring denser configurations in high-compute regimes.
We release model checkpoints, code and logs are open-source at https://github.com/rioyokotalab/optimal-sparsity.
@@ 2.1 Mixture of Experts
Mixture-of-Experts (MoE) networks were introduced by (Jacobs et al., 1991; Jordan & Jacobs, 1994) and later brought to large-scale neural language modeling by Shazeer et al. (2017). Within the Transformer architecture (Vaswani et al., 2017), MoE layers have proven especially effective, scaling to hundreds of billions of parameters while maintaining manageable training costs (Lepikhin et al., 2021; Fedus et al., 2021; Du et al., 2022). In an MoE layer, a learnable router assigns each token to a sparse subset of experts. Let $\mathbf{x}\in\mathbb{R}^{d_{h}}$ be a token representation and $\{\text{FFN}(\mathbf{x})_{i}\}_{i=1}^{n}$ the $n$ feed-forward experts. For top-$k$ routing, the router produces scores $\mathbf{s}=\mathbf{x}^{\top}\mathbf{W}_{\text{router}}\in\mathbb{R}^{n},$ and selects the indices $\mathcal{K}$ of the $k$ largest components, then normalizes them: $g(\mathbf{x})_{i}=\frac{\exp(s_{i})}{\sum_{j\in\mathcal{K}}\exp(s_{j})}$ if $i\in\mathcal{K}$ and $g(\mathbf{x})_{i}=0$ otherwise.
@@ 2.2 Scaling Laws of LLMs
Existing scaling laws demonstrate power-law relationships between model performance, parameter count, dataset size, and compute budget (Kaplan et al., 2020; Hoffmann et al., 2022). Scaling laws for MoE models have similarly explored how total parameter count and expert granularity jointly affect scaling behavior (Clark et al., 2022; Ludziejewski et al., 2024). Building on this, Frantar et al. (2024) derived sparsity-aware scaling exponents that bridge dense and sparse regimes, while Abnar et al. (2025) empirically charted the optimal trade-offs between total parameters and FLOPs per token in MoE settings. Furthermore, recent analyses indicate that increasing sparsity itself can directly improve loss, highlighting sparsity as a key dimension in scaling behavior (Kimi Team, 2025).
@@
Since the scaling law for next-token prediction loss does not necessarily align with downstream task loss, it may not be reliable for predicting benchmark performance (Grattafiori et al., 2024). Some work has tried to model downstream accuracy with an exponential curve, but accuracy is only predictable when we average over many tasks and carefully choose which ones to include (Gadre et al., 2024). Another line of research instead first quantifies how downstream task loss scales with parameters and data, then converts predicted losses into accuracy estimates, achieving under two points of absolute error for mid-scale models using minimal extra compute (Bhagia et al., 2024). Prior work observes that downstream task loss relates to pre-training loss, where the shifts depend on the minimal achievable losses determined by the intrinsic complexity and distributional mismatch between the pre-training and downstream datasets (Brandfonbrener et al., 2025).
Skills
Adding MoE experts tends to improve memorization skills more than reasoning skills, motivating new, generalized scaling frameworks that address scaling laws for reasoning performance (Jelassi et al., 2025). scaling laws differ across tasks, the optimal scaling strategy may also vary; for example, knowledge-based QA tasks are “capacity-hungry,” benefiting more from larger model sizes, whereas code-related tasks are “data-hungry,” benefiting more from increased training data (Roberts et al., 2025). Complementing these observations, recent analyses find that models with identical training loss can still exhibit markedly different reasoning performance (GLM-4.5 Team, 2025), highlighting that reasoning skill depends not only on loss but also on how compute and data are allocated.
@@ 3 Experiments
In this section, we empirically demonstrate the scaling of downstream task performance through a systematic investigation of memorization and reasoning skills in MoE LLMs.
@@ 3.1 Experimental Setup
We use a balanced 125B-token mixture consisting of high-quality web text (43B), mathematics corpora (32B), STEM literature and reference (49B), and code (1B). See Appendix A.1 for complete statistics.
@@
We evaluate three capability areas with standard few-shot prompts. Mathematical Reasoning: GSM8K (Cobbe et al., 2021) (4-shot) and GSM-Plus (Li et al., 2024) (5-shot CoT). Reading Comprehension: TriviaQA (Joshi et al., 2017) with 4-shot prompting. Commonsense Reasoning: HellaSwag (Zellers et al., 2019), each under a 4-shot prompting setup. See Table 3 in Appendix for further details.
@@ 3.2 Downstream Performance Does Not Necessarily Improve with Total Parameter Size
Figure 1 presents the training and validation losses when fixing the top-$k$, MoE layer width constant and increasing only the number of experts (and hence the total parameter count). As the total parameter count grows, both training and validation losses decrease. Therefore, in terms of pre-training loss, increasing total parameters (thereby raising sparsity) reduces pre-training loss, which is consistent with prior work.
@@
Can we avoid a decline in performance as the total number of experts increases? Figure 2 shows that models with more active parameters begin to overfit at a lower pre-training loss and reach a lower minimum task loss at their optimal expert counts. Consequently, improving results on GSM8K and GSM-Plus requires tuning not only the total number of experts but also the top-$k$ size.
@@
The decline in math-task performance as total parameters increase is not limited to task loss; it also consistently holds for downstream accuracy (Figure 3). For TriviaQA and HellaSwag, accuracy improves monotonically as training loss decreases. By contrast, on GSM8K, further reductions in pre-training loss do not always translate to higher accuracy. When the number of active parameters is held constant, over-optimizing pre-training loss can indeed harm performance. Figure 4 plots benchmark error rate against pre-training loss, including intermediate checkpoints. We observe a sparsity dependence for reasoning skills. These results suggest that, for MoE models, downstream accuracy can deviate from the predictions of conventional scaling laws, and these deviations may vary across different tasks. This effect persists even when controlling for optimization hyperparameters such as learning rate (Appendix C.7).
@@
Figure 4: Effect of sparsity on performance across different tasks We vary sparsity (1 - top-kk/Experts) and plot the relationship between pre-training loss and benchmark error rate, including intermediate checkpoints. For TriviaQA and HellaSwag, the error rate clearly tracks training loss and is largely insensitive to sparsity. In contrast, reasoning skills exhibit a strong dependence of error rate on sparsity.